OpenStudy (anonymous):

A particle has: A.) an average acceleration of 6 m/s2 if it goes from 2 m/s to 6 m/s in 2 seconds. B.) an average acceleration of 4 m/s2 if it goes from 2 m/s to 6 m/s in 2 seconds. C.)an average acceleration of 12 m/s2 if it goes from 2 m/s to 6 m/s in 2 seconds. D.)an average acceleration of 3 m/s2 if it goes from 2 m/s to 6 m/s in 2 seconds. E.) an average acceleration of 2 m/s2 if it goes from 2 m/s to 6 m/s in 2 seconds.

OpenStudy (ash2326):

@mikethorp Do you know the formula relating acceleration and velocity?

OpenStudy (ash2326):

It's \[a= \frac{v-u}{t}\] a= acceleration in m/s^2 v= final velocity in m/s u=initial velocity in m/s t= time in seconds

OpenStudy (ash2326):

Can you use this to find the correct option? @mikethorp

OpenStudy (anonymous):

Yea